OpenStudy (lovelycharm):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 16 y = x + 4

Vocaloid (vocaloid):

do a substitution y = x + 4 means we can replace "y" with "x + 4" when we see y, so x^2 + y^2 = 16 x^2 + (x+4)^2 = 16 know where to go from here?

OpenStudy (lovelycharm):

then u divide the by 4

Vocaloid (vocaloid):

uhhhh not quite

Vocaloid (vocaloid):

then we can use FOIL to expand (x+4)^2

OpenStudy (lovelycharm):

2x+8

Vocaloid (vocaloid):

not quite, the ^2 is an exponent (x+4)^2 = (x+4)(x+4) = ?

OpenStudy (lovelycharm):

0

Vocaloid (vocaloid):

nope, not quite... hint: (x+A)(x+A) = x^2 + 2*A*x + A^2 (x+4)(x+4) = ____________________ fill in the blank

OpenStudy (lovelycharm):

x+4x+4x+16

Vocaloid (vocaloid):

almost, you're very close! x^2 + 4x + 4x + 16

Vocaloid (vocaloid):

that simplifies to: x^2 + 8x + 16, with me so far?

OpenStudy (lovelycharm):

2x+8x+16 4x+16

Vocaloid (vocaloid):

nope, not quite. please remember that x^2 is NOT equal to 2x

Vocaloid (vocaloid):

x^2 means x*x, ok?

OpenStudy (lovelycharm):

so it was x+8x+16

OpenStudy (lovelycharm):

confuse

Vocaloid (vocaloid):

yes, x^2 + 8x + 16, with me so far?

OpenStudy (lovelycharm):

x+4^2

Vocaloid (vocaloid):

no, we just leave it as x^2 + 8x + 16, ok?

OpenStudy (lovelycharm):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 16 y = x + 4 −4 and 0 4 and 0 −2 and −3 2 and −3

OpenStudy (lovelycharm):

those r the option

Vocaloid (vocaloid):

right, we'll get to that in a minute, ok?

Vocaloid (vocaloid):

right now we have: (x+4)^2 = x^2 + 8x + 16 plugging that back into the first equation gives us x^2 + x^2 + 8x + 16 = 16 know where to go from here?

OpenStudy (lovelycharm):

no not really

Vocaloid (vocaloid):

start by subtracting 16 from each side.

OpenStudy (lovelycharm):

0

Vocaloid (vocaloid):

what about the left side? x^2 + x^2 + 8x + 16 - 16 = ?

OpenStudy (lovelycharm):

4

Vocaloid (vocaloid):

nope, maybe this will make things easier: (x^2 + x^2 + 8x) + 16 - 16 = ?

OpenStudy (lovelycharm):

8x^2+0

OpenStudy (lovelycharm):

2x^2+8x

Vocaloid (vocaloid):

good, 2x^2 + 8x

Vocaloid (vocaloid):

now we have 2x^2 + 8x = 0 solve for x

OpenStudy (lovelycharm):

4x+8x=0 4x=0

Vocaloid (vocaloid):

remember what I said earlier? x^2 is NOT 2x

Vocaloid (vocaloid):

2x^2 + 8x = 0 2x(x+4) = 0 solve for x

OpenStudy (lovelycharm):

2x2+8x

Vocaloid (vocaloid):

not quite... 2x^2 + 8x is what I gave you to start with

Vocaloid (vocaloid):

2x = 0 x = ?

OpenStudy (lovelycharm):

4

Vocaloid (vocaloid):

try again

Vocaloid (vocaloid):

2x = 0 divide both sides by 2

OpenStudy (lovelycharm):

0

Vocaloid (vocaloid):

good

Vocaloid (vocaloid):

now we solve (x+4) = 0 x = ?

OpenStudy (lovelycharm):

4

Vocaloid (vocaloid):

almost but not quite

Vocaloid (vocaloid):

subtract 4 from each side

OpenStudy (lovelycharm):

-4

Vocaloid (vocaloid):

good, so our solutions are 0 and -4

OpenStudy (lovelycharm):

thank u so much